OpenStudy (anonymous):

Help with these triangle problems? Thanks in advance!! 1. A triangle has angle measure of 40°, 59°, and 81°. The side opposite of the 40° angle is 11cm long. Find the lengths of the side opposite the 81° angle to the nearest centimeter. 2. Given triangle ABC with a=14, A=25°, B=17, find c. Round to the nearest whole number. 3. Solve triangle ABC given that a=12, b=19, c=21. Round to the nearest degree. Thanks, again.

OpenStudy (zzr0ck3r):

do you know about law of sines?

OpenStudy (anonymous):

@zzr0ck3r No :(

OpenStudy (anonymous):

Is there a formula for the law of sines?

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (zzr0ck3r):

sorry, um yeah but if you are not supposed to use it, then we should use some other method...what are you doing in class?

OpenStudy (anonymous):

This is actually an online class

OpenStudy (zzr0ck3r):

what are you doing in the online class?

OpenStudy (zzr0ck3r):

if you want to use law of sines then here is how it works  sin(a)/A = sin(b)/B = son(c)/C

OpenStudy (zzr0ck3r):

you have a = 40 and A = 11, and b = 81 so sin(40)/11 = sin(81)/x solve for x x=11*sin(81)/sin(40)